OpenStudy (anonymous):

what is the y-intercept of the line? y=-3x

OpenStudy (studio-games):

Y intercept is -9

OpenStudy (sleepyhead314):

The y intercept is when x = 0 so try plugging in 0 for x and see what you get :) y = -3*(0) = ?

Vocaloid (vocaloid):

hint: y = -3x + 0 y = mx + b b = y-intercept = 0

OpenStudy (anonymous):

i have A.-3 B.0 C.1 D.3

OpenStudy (sleepyhead314):

@studio-games Hi :) I appreciate your effort there, but it would really help if you can include an explanation with how you got to your answer ^_^

OpenStudy (sleepyhead314):

@austinhug Thanks for including the options, try taking a look at the hint that @Vocaloid gave :)

OpenStudy (anonymous):

it doesn't make seance to me

OpenStudy (foreverandalways13):

okay try this the orginal equation is y=mx+b and your life is easy b=0 because b doesnt exist in this problem

OpenStudy (foreverandalways13):

does that help?

OpenStudy (foreverandalways13):

b is the y intercept but in the equation there isnt a b

OpenStudy (foreverandalways13):

an example with a b would be y=3x+2